• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name pking

    Log entry time 00:30:33 on September26,2005

    Entry number 155966

    keyword=Swing shift summary

    16:00---Begin shift with parity run 5365 and Compton run 12298 in progress.
    
    16:57---Begin parity run 5366, and Compton run 12299.
    
    17:35---Bryan is trying to reboot the spectrometer DAQ ROCs.  They seem
    	to be having trouble connecting to the network.
    
    17:54---End parity run 5366, and Compton run 12299.
    	
    Bryan begins to set up the position feedback with the helicity magnets.  The 
    tests of position feedback using the helicity magnets are described in HALog 
    entry 155942
    
    17:54---Begin Compton run 12300.
    
    17:58---The IHWP is inserted.
    
    19:15---Found Compton run 12300 with DAQ problem; ROC0 was not responding.  
    Reboot CODA.  CODA would not start properly.  Reboot CODA again.  CODA starts, 
    and begin run 12301.
    
    
    20:13---Begin run 5374.  The helicity magnets are off.
    
    20:25---End of the helicity magnet tests; Bryan will analyze the data offline 
    to figure out why the feedback did not work.
    	Begin parity run 5375 and Compton run 12302.
    
    21:20---Compton run 12302 crashed.  Begin Compton run 12303, begin parity run 
    5376.
    
    
    21:45---Moller quad trip, beam will be off ~15 min
    
    21:54---They can not turn on the Moller quads.  Minimum estimate would be an 
    hour before the quads would be fixed (0.5 hr for the personel to arrive, 0.5 hr 
    to fix it)
    
    22:23---The RC said that we shouldn't run without the Moller quads, especially 
    since the spectrometer DAQ and spot++ aren't working and we can't check the 
    beam spot size.
    
    22:34---Move the target to the empty position.
    
    22:40---The left HRS dipole has tripped.
    
    22:51---Bryan is going into the hall with the ARM to try to get the 
    spectrometer DAQ working.  He has rebooted the network switch.
    
    22:58---Paged Tech-On-Call about the Left Dipole.  His instructions are to turn 
    off the power supply for two minutes from the switch on the front panel.
    
    23:45---The left dipole has been reset and is now at its nominal current.  
    Bryan remains in the hall to begin opening the spectrometer huts to allow work 
    on the profile scanners.
    
    
    Parity run summary:
    
    Run 5365: IHWP OUT   48k pairs
    Run 5366: IHWP OUT   49k pairs
    Run 5375: IHWP IN    47k pairs
    Run 5376: IHWP IN    18k pairs
    
    Compton run summary:
    Run 12298: IHWP OUT  (81.6 +/- 0.8 +/- 3.7)%
    Run 12299: IHWP OUT  (83.6 +/- 1.1 +/- 3.7)%
    Run 12302: IHWP IN   (-83.3 +/- 1.1 +/- 3.7)%
    Run 12303: IHWP IN   (-85.2 +/- 1.7 +/- 3.7)%